# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

### v1

@@ 3 Strongly Convex Setting
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq{a_{0}\|x_{0}-x_{*}\|^{2}}+{e^{\text{sum}}\sigma_{*}^{2}},\quad\text{ with }\quad e^{\text{sum}}=\sum_{t=0}^{T-1}e_{t}.$
We now present our main results for the strongly convex case, which are obtained after carefully choosing Lyapunov parameters in Appendix D. The first theorem provides a bound with sharp bias terms for all step-sizes, except for the so-called optimal step-size $\gamma=\frac{2}{L+\mu}$ for which the variance explodes. The second theorem provides a bound valid for every step-size (including the optimal one), with an $\varepsilon$-sub-optimal bias term which allows for a finite variance term.
Theorem 3.3 (Strongly convex case, sharp bias).
Let Assumption 3.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,2)$ such that $\gamma\neq\gamma_{opt}=\tfrac{2}{\mu+L}$. Then, for every $T\geq 1$,
where $\phi=\max\{1-\gamma\mu;\gamma L-1\}\in[0,1)$, and $e=\gamma^{2}\left(1+\tfrac{\gamma(L-\mu)}{|\gamma-\gamma_{opt}|(L+\mu)}\right)$.
Theorem 3.4 (Strongly convex case, sub-optimal bias).
Let Assumption 3.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,2)$. Then, for every $T\geq 1$ and $\varepsilon>0$,
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e=\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon}\right)$.
As in the convex case, when interpolation holds, namely when $\sigma_{*}^{2}=0$, the division by $\varepsilon$ is no longer a problem, allowing us to take $\varepsilon\to 0$.
Corollary 3.5 (Strongly convex case with interpolation, optimal bias).
Let Assumption 3.1 hold, and assume that $\sigma_{*}^{2}=0$. Let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,2)$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2},$
where $\phi=\max\{1-\gamma\mu;\gamma L-1\}\in[0,1)$.
The first message to take from our results is that it is possible to have bounds for SGD whose bias term corresponds exactly to the standard rate for GD, namely a geometrical rate governed by $\phi^{2}$. A second insight is that it is possible to obtain bounds for the full range of step-sizes $\gamma L\in(0,2)$. This is in contrast with (bach_non-asymptotic_2011, Theorem 1), which employs a similar Lyapunov approach to ours, and obtains a bias term governed by $\phi^{2}+\gamma^{2}(L^{2}-\mu^{2})$ for step-sizes $\gamma L\in(0,\tfrac{\mu}{L})$. We can also compare with (needell_stochastic_2016, Theorem 2.1), which provides an improved bias term governed by $\phi^{2}+2\gamma^{2}\mu(L-\mu)$ for short step-sizes $\gamma L\in(0,1)$, or the more recent (gower_sgd_2019, Theorem 3.1), whose bias term is governed by $\phi$ (with no square) for step-sizes $\gamma L\in(0,\tfrac{1}{2})$.
It is remarkable that being able to work with large step-sizes allows us to observe a singular phenomenon for the optimal step-size, which, to our knowledge, was never observed before. In the deterministic setting, the step-size $\gamma=\tfrac{2}{\mu+L}$ is the only one leading to the optimal geometric rates governed by $\phi^{2}=\tfrac{(L-\mu)^{2}}{(L+\mu)^{2}}$. In our stochastic setting, it appears that achieving such a bias term is incompatible with having a finite variance. It seems reasonable to think that achieving the optimal rate for a deterministic method requires such precision that introducing a small perturbation or variance might break it. Such conjecture seems to be validated by our numerical experiments presented in the next section.
4 Sharpness of our bounds
In this section, we delve into discussing the tightness of our bounds. To this end, we employ a numerical approach based on the Performance Estimation Problem (PEP) methodology, initially introduced by drori_performance_2014, further improved by taylor_smooth_2017, and adapted by taylor_stochastic_2019 in the context of stochastic algorithms, that we will follow closely. We direct the reader to Appendix E for an in-depth explanation of our PEP framework. Roughly speaking, the idea consists in reformulating the problem of finding admissible Lyapunov parameters ensuring the energy decrease into a semi-definite program, which may be solved numerically. For this paper, we focused on finding admissible coefficients yielding minimal bias terms in our bounds.
This framework inspired and drove our proofs, helping us to postulate the best Lyapunov parameters that could be found. We made this clear in our proofs, which can be found in the Appendices C and D. Although most of our theorems in Sections 2 and 3 are stated using simplified values, for an easier presentation, in this section we will make use of the exact values found in our proofs. Moreover, it allowed us to verify the tightness of our bounds, at least numerically. We recognize that theoretical results overlapping with numerical observations do not constitute a proof of optimality. Nevertheless, these validations allow us to formulate conjectures that our results are sharp, some of which we partially prove. This tightness assessment will be the focus of the rest of this section.
All of our numerical experiments [^1] Code provided in supplementary material. We will provide a link to a public repository in the final version. were developed in Python 3.13 using the solvers MOSEK version 11.0.19 (mosek_mosek_2025) and Clarabel version 0.9.0 (goulart_clarabel_2024) to solve the resulting semi-definite programs, implemented using CVXPY version 1.6.5 (diamond_cvxpy_2016), and run on Intel Xeon Platinum 8380 CPUs.
4.1 Convex Setting
The primary focus of our work has been to identify sharp bias terms, which directly translates to maximizing the value of $\rho$ via Lemma 2.2, once we assume, without loss of generality, that $a_{0}=1$. We denote by $\rho_{\text{opt}}$ the supremum of all such possible admissible constants $\rho$. In parallel, we define $\rho_{\text{theory}}$ to be our guess for $\rho_{\text{opt}}$ driven by our results in Theorems 2.3, 2.4 and 2.7:
$\rho_{\text{theory}}\coloneqq\begin{cases}2\gamma+\tfrac{2}{TL}(1-\gamma L)&\text{ if }\gamma L\in(0,1),\\ 2\gamma&\text{ if }\gamma L=1,\\ 2\gamma(2-\gamma L)\left[1-(1-\gamma L)^{2T}\right]^{-1}&\text{ if }\gamma L\in(1,2).\end{cases}$
Figure 1 compares our analytically derived $\rho_{\text{theory}}$ with the value of $\rho_{\text{opt}}$ obtained numerically. The perfect match across the full range of step-sizes validates the sharpness of the bias term in our theoretical bounds for step-sizes $\gamma L\neq 1$. This empirical sharpness is validated by a formal proof, for large step-sizes (see Proposition C.10 for more details).
Proposition 4.1.
For large step-sizes $\gamma L\in(1,2)$, there exists no Lyapunov parameters with $a_{0}=1$ and $\rho>\rho_{\text{theory}}$ making the energy decrease.
For the optimal step-size $\gamma L=1$, we empirically see that we cannot do better than $2\gamma$, but we were not able to achieve it in Theorem 2.4. To understand why, we turn to the study of the variance term.
Figure 1: Theoretical and numerical bias term ($L=1$).
As a second set of experiments, we minimized the constant $\overline{e}$ (which governs the variance term in our bound from Lemma 2.2) under the constraint that $\rho=\rho_{\text{theory}}$. We will denote the infimum of such constants as $\overline{e}_{\text{opt}}$, and will note $\overline{e}_{\text{theory}}$ the constant we obtained in our results. We compare the values of those two constants in Figure 2, for non-optimal step-sizes. We can see that, as $\gamma L\to 1$, either from the left or from the right, the variance term tends to infinity, suggesting that it is not possible to maintain $\rho=\rho_{\text{opt}}$ when $\gamma L=1$. Moreover, the solver fails when trying to compute $\overline{e}_{\text{opt}}$ for this optimal step-size, which again points to the fact that $\rho=2\gamma$ is not a feasible Lyapunov parameter when $\gamma L=1$. This is why in Theorem 2.4 we proposed a bias term with any $\rho_{\varepsilon}<\rho_{\text{opt}}$, in view of maintaining a finite variance. In consequence we conjecture that, within such Lyapunov framework, it is not possible to obtain a bound for SGD with a bias term of the order $\tfrac{1}{2\gamma T}$ when $\gamma L=1$.
A bi-product of the results displayed in Figure 2 is that the variance terms we obtained in Theorems 2.3 and 2.7 overlap with the variance returned by the solver. This suggests that we cannot avoid the issue we had with Theorem 2.7, where the variance term grows exponentially with time.
We note that the problem, specially for large step-sizes, is highly numerically unstable, which is why we do not provide plots for $T>2$. This is further discussed in Appendix F.
Figure 2: Theoretical and numerical average variance term ($L=1$).
4.2 Strongly Convex Setting
In this setting, our focus has also been to identify sharp bias terms which, through Lemma 3.2, amounts to minimizing the value of $a_{0}$, once we assume, without loss of generality, that $a_{T}=1$. We denote by $a_{0}^{\text{opt}}$ the infimum of all such possible admissible constants. Driven by our result in Theorem 3.3, we expect $a_{0}^{\text{opt}}$ to be equal to $\phi^{2T}$, with $\phi=\max\{1-\gamma\mu,\gamma L-1\}$. Figure 3 compares $\phi^{2T}$ with the value of $a_{0}^{\text{opt}}$ obtained numerically, and illustrates again the tightness of our results. Note that the iteration count is smaller here than in the convex setting, due to numerical instability issues, which we detail in Appendix F.
Figure 3: Theoretical and numerical bias term ($L=1$, $\mu=0.25$ and $\tfrac{2}{\mu+L}=1.6$).
For the optimal step-size $\gamma=\frac{2}{L+\mu}$, we empirically see that we cannot do better than $\phi^{2T}$, but we were not able to achieve it in Theorem 3.3. Here again we will understand why by looking at the variance terms. To do so, we implement the minimization of the constant $e^{sum}$ appearing in Lemma 3.2 under the constraint that $a_{0}$ is equal to $\phi^{2T}$. Figure 4 compares the optimal value $e^{sum}_{\text{opt}}$ to the constant we obtained in our proofs. We can see that, as the step-size gets closer to the optimal one, the variance tends to +$\infty$+ ∞. This strongly suggests that it is not possible to obtain a bias term of the order $\left(\tfrac{L-\mu}{L+\mu}\right)^{2T}$ by means of a Lyapunov analysis such as ours.
Figure 4: Theoretical and numerical total variance term ($L=1$, $\mu=0.25$ and $\tfrac{2}{\mu+L}=1.6$).
5 Conclusion and perspectives
In this paper we proved new and improved upper bounds for SGD without any variance assumption, both in the convex and strongly convex setting, for a full range of step-sizes. We moreover illustrated the sharpness of these results numerically using the Performance Estimation Problem methodology. Doing so, we raised questions regarding unexpected singularities appearing for optimal step-sizes.
5.1 Extensions
For the sake of presentation, we formulated our main results under strong but easy to grasp assumptions, and left the more general and technical assumptions to Appendix A. For instance, all of our results remain valid when replacing the Problem (Finite-Sum) with the minimization of an expectation $\mathbb{E}\mathopen{}\left[{f_{i}}\right]\mathclose{}$ of a (possibly infinite) family of functions $f_{i}$. Moreover, Assumptions 2.1 and 3.1 may be replaced by a weaker symmetric expected cocoercivity assumption.
We moreover note that the analysis carried out for (SGD) extends to the case of mini-batching and to non-uniform sampling without much additional effort. This is explained in Appendix G.
5.2 Future Work
We leave the following questions open and lines of research:
  • Are the upper bounds obtained improvable through more general forms of Lyapunov energies? We worked with an energy of the form (Lyapunov), but do not claim this is the only sensible form.
  • Is it possible to obtain upper bounds on (SGD) whose bias term match the best upper bound in the deterministic gradient descent algorithm? Both in the convex and strongly convex setting, we still observe a gap, and we conjecture this is not due to non-tightness of our analysis but rather a flaw in the core analysis technique.
  • The analysis carried out solely focuses on constant step-sizes. Can this type of analysis be extended to cover varying step-sizes?
  • In this paper, we focused on achieving sharp bias terms. However, it is not the only possible approach, and we intend in a future work to investigate bounds providing sharp complexity rates.
Acknowledgments and Disclosure of Funding
This work benefited from the support of the FMJH Program Gaspard Monge for optimization and operations research and their interactions with data science. We also thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high performance computing cluster.
Appendix A Problem definition and main assumptions
Throughout the appendix sections, we consider a more general setting than in Sections 2 and 3. Instead of limiting ourselves to the minimization of a finite sum of functions in $\mathbb{R}^{d}$, we consider the minimization of an expectation of functions defined on a real Hilbert space.
Problem A.1.
[…]
@@ Appendix A Problem definition and main assumptions
Within the context of Problem A.1, we study the performance of the SGD algorithm, which generates a sequence $(x_{t})_{t\geq 0}$ according to
$x_{t+1}=x_{t}-\gamma\nabla f_{i_{t}}(x_{t}),$ (SGDE)
where $\gamma>0$ is a fixed step-size, and $i_{t}\in\mathcal{I}$ is sampled i.i.d. from the distribution $\mathcal{D}$.
A.1 Smoothness and convexity
The main assumptions that we present in the core of the paper are smoothness and convexity (or strong convexity) of the functions $f_{i}$. We define a setting combining both convexity and strong convexity, which allows us to unify parts of the analyses. We also define a more relaxed setting, which is all that is needed for our results to hold.
Assumption A.2 (Smoothness and $\mu$-strong convexity).
Considering Problem A.1, we assume that
@@ A.1 Smoothness and convexity
  • there exists $L\in(0,+\infty)$ such that each function $f_{i}$ is $L$-smooth, in the sense that $\nabla f_{i}\colon\mathcal{H}\to\mathcal{H}$ is $L$-Lipschitz continuous;
@@
  • there exists $\mu\in[0,+\infty)$ such that each function $f_{i}$ is $\mu$-strongly convex. When $\mu=0$, this simply means that $f_{i}$ is convex.
@@
Smoothness and $\mu$-strong convexity can be characterized by the means of a variational inequality:
Lemma A.3 (Expected Cocoercivity).
Let Assumption A.2 hold. Then, for every $(x,y)\in\mathcal{H}^{2}$,
@@
$\ \frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(y)-\nabla f_{i}(x)\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|y-x\|^{2}$
$\leq$ $\ (L-\mu)\left(f(y)-f(x)\right)+\langle\mu\nabla f(y)-L\nabla f(x),y-x\rangle.$ (EC)
Proof.
@@
Because $f_{i}$ is $L$-smooth and $\mu$-convex, we can use (taylor_smooth_2017, Theorem 4) to write, at every $(x,y)$:
$\frac{1}{2}\|\nabla f_{i}(y)-\nabla f_{i}(x)\|^{2}+\frac{\mu L}{2}\|y-x\|^{2}$
$\leq$ $(L-\mu)\left(f_{i}(y)-f_{i}(x)\right)+\langle\mu\nabla f_{i}(y)-L\nabla f_{i}(x),y-x\rangle.$
The conclusion follows after taking expectation. ∎
As will be made clear throughout our proofs, we may relax Assumption A.2 to only assume that the conclusion of Lemma A.3, namely (EC), holds between any point $x$ and any minimizer $x^{*}$. We will decline this assumption into two flavors.
Assumption A.4 (Expected Cocoercivity).
Considering Problem A.1, we assume that there exist $L\in(0,+\infty)$ and $\mu\in[0,L]$ such that
for every $x\in\mathcal{H}$ and every $x_{*}\in{\rm{argmin}}~{}f$, (EC) holds for $(x,x_{*})$. (EC)
Assumption A.5 (Symmetric Expected Cocoercivity).
Considering Problem A.1, we assume that there exist $L\in(0,+\infty)$ and $\mu\in[0,L]$ such that
for every $x\in\mathcal{H}$ and every $x_{*}\in{\rm{argmin}}~{}f$, (EC) holds for $(x,x_{*})$ and $(x_{*},x)$. (SEC)
Remark A.6 (Comparing (SEC) and (EC) with other assumptions in the literature).
From what precedes, it is clear that
each $f_{i}$ is $\mu$-convex and $L$-smooth $\Rightarrow$ (SEC) $\Rightarrow$ (EC).
We also note that (EC) implies the $\mu$-strongly quasi-convexity assumption from gower_sgd_2019. If we focus on the case $\mu=0$, we see that (SEC) holds if, and only if, (EC) and (ES) hold, where the latter is the Expected Smoothness property introduced by gower_sgd_2019, stating that, for all $x\in\mathcal{H}$ and $x_{*}\in{\rm{argmin}}~{}f$,
$\frac{1}{2L}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})-\nabla f_{i}(x)\|^{2}}\right]\mathclose{}\leq f(x)-f(x_{*}).$ (ES)
A.2 Gradient variance at the solution
We also assume that the variance of $\nabla f_{i}(x)$ exists at the solutions. More precisely,
Assumption A.7 (Solution Gradient Variance).
Considering Problem A.1, we assume that the variance at the solution exists, meaning that
$\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]<+\infty$ (GV)
for every $x_{*}\in{\rm{argmin}}~{}f$. We will note
@@ A.2 Gradient variance at the solution
$\sigma_{*}^{2}\coloneqq\sup\limits_{x_{*}\in{\rm{argmin}}~{}f}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}.$
@@
It is important to note that under Assumption A.4 the variance term $\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ does not depend on the choice of $x_{*}\in\operatorname{argmin}f$, as was already observed in (garrigos_handbook_2024, Lemma 4.17).
Lemma A.8.
@@
Let Assumption A.4 hold. Then $\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ is constant over ${\rm{argmin}}~{}f$.
Proof.
@@
Let $x_{*},x_{*}^{\prime}$ be two minimizers of $f$. From Assumption A.4 we know that (EC) holds true for the couple $(x_{*},x_{*}^{\prime})$, which means that
@@
$\frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$
$\leq$ $(L-\mu)\left(f(x_{*}^{\prime})-f(x_{*})\right)+\langle\mu\nabla f(x_{*}^{\prime})-L\nabla f(x_{*}^{\prime}),y-x\rangle.$
@@
Then, almost surely with respect to the distribution $\mathcal{D}$ over the indices $i\in\mathcal{I}$, we have $\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}=0$. This means that $\nabla f_{i}(x_{*}^{\prime})=\nabla f_{i}(x_{*})$ almost surely, from which we deduce that
$\mathbb{E}\left[\|\nabla f_{i}(x_{*}^{\prime})\|^{2}\right]=\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right],$
meaning that in fact $\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ is constant over ${\rm{argmin}}~{}f$. ∎
Assuming that the variance at the solution $\sigma_{*}^{2}$ exists is trivially satisfied in the case of the minimization of a finite sum, as presented in Problem (Finite-Sum).
Lemma A.9 (Finite-sum problems have finite solution variance).
If Problem A.1 treats a finite sum of functions, i.e. $\mathcal{I}$ is finite, then Assumption A.7 holds true.
Even for true expectation-minimization problems, where $\mathcal{I}$ is infinite, the Assumption A.7 is very mild. In particular, this assumption is automatically verified in problems where the loss functions $f_{i}$ are nonnegative, which is standard for problems arising in inverse problems and machine learning.
Lemma A.10 (Sufficient condition for finite solution variance).
Consider Problem A.1 and assume that the functions $f_{i}$ are $L$-smooth and are uniformly bounded from below functions: there exists $r\in\mathbb{R}$ such that $f_{i}(x)\geq r$ for every $i\in\mathcal{I}$ and $x\in\mathcal{H}$. Then Assumption A.7 is true, with moreover the information that
@@
Proof.
Since each $f_{i}$ is $L$-smooth, we can write (garrigos_handbook_2024, Lemma 2.28), for all $x\in\mathcal{H}$,
@@
$\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]\leq 2L(f(x_{*})-r)=2L(\min f-r)<+\infty,$
which concludes since $\sigma_{*}^{2}=\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$. ∎
A.3 Interpolation
We conclude by connecting the variance at the solution with interpolation. The following facts can be found in garrigos_handbook_2024, which we have extended here from a finite-sum problem to our expectation-minimization Problem A.1.
Definition A.11.
Consider Problem A.1. We say that interpolation holds for the family of functions $(f_{i})_{i\in\mathcal{I}}$ if
$\bigcap\limits_{i\in\mathcal{I}}\ {\rm{argmin}}~{}f_{i}\neq\emptyset,$
where the intersection is to be understood for $\mathcal{D}$-almost every $i\in\mathcal{I}$.
Lemma A.12 (Consequences for interpolation).
Consider Problem A.1. If interpolation holds, then
  1. 1.
    $\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}=\inf f$.
  2. 2.
    $\cap_{i\in\mathcal{I}}\ {\rm{argmin}}~{}f_{i}={\rm{argmin}}~{}f$.
Proof.
@@ A.3 Interpolation
Since $f_{i}(x)-\inf f_{i}\geq 0$ for all $i$, it must hold that $f_{i}(x)-\inf f_{i}=0$ for almost every $i\in\mathcal{I}$, which proves the claim. ∎
Lemma A.13 (Interpolation means $\sigma_{*}^{2}=0$).
Consider Problem A.1.
  1. 1.
    If interpolation holds, then $\sigma_{*}^{2}=0$.
  2. 2.
    The above becomes an equivalence if all the $f_{i}$ are convex.
Proof.

### v2

@@ 1 Introduction
$\sup_{x\in\mathcal{H}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)-\nabla f(x)\|^{2}}\right]\mathclose{}<+\infty\quad\text{or}\quad\sup_{x\in\mathcal{H}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)\|^{2}}\right]\mathclose{}<+\infty.$
Such assumptions date back to the early works of the 1950s (chung_stochastic_1954, ; sacks_asymptotic_1958, ) and continued to be employed until recently (polyak_acceleration_1992a, ; nemirovski_robust_2009, ). A culminating contribution to the nonasymptotic analysis of SGD under bounded variance assumptions is taylor_stochastic_2019 , where new and sharp convergence rates were established. Over time, these variance assumptions have been progressively relaxed and replaced by conditions that control the gradient or its variance through quantities such as $\|x\|$ or $\|\nabla f(x)\|$, we refer to the overview in alacaoglu_towards_2025 and the references therein. Nevertheless, variance-based assumptions can be unrealistic in practice (bottou_optimization_2018, ; nguyen_sgd_2018, ). Moreover, uniform boundedness of the gradients can never be satisfied for strongly convex functions.
A new line of research was initiated in bach_non-asymptotic_2011 by removing any such variance assumptions, followed by the more recent works (needell_stochastic_2016, ; nguyen_sgd_2018, ; gower_sgd_2019, ; gower_sgd_2021, ; khaled_better_2023, ; demidovich_guide_2023, ). The core idea is to exploit the convexity and smoothness to obtain a variance transfer inequality, bounding the variance at any point with a term involving $\sigma^{2}_{*}$. The new gold standard for the analysis of SGD quickly became to not make any variance assumption when assuming convexity and smoothness, a framework into which our work fits.
We point out that convexity and smoothness is a sufficient but not a necessary condition to obtain complexity results: all one needs is for some variance transfer inequality to hold. We mention for instance the expected smoothness assumption (gower_stochastic_2021, ; gower_sgd_2019, ; khaled_better_2023, ) and the ABC condition (demidovich_guide_2023, ), which were shown to be sufficient to establish performance guarantees. These assumptions are strictly weaker than convexity and smoothness, enabling the analysis of SGD in a nonconvex framework. While our analysis could also be carried out under weaker assumptions than convexity and smoothness, it is unclear how practical such assumptions would be, and we do not delve further into this in this work.
Note on Concurrent Work. In the convex setting, our work focuses on rates on the ergodic function value gap $f(\bar{x}_{T})-\min f$, which is currently the standard for SGD without uniformly bounded variance. In contrast, when making a stronger assumption on the stochastic gradients, it is standard to obtain rates on the last-iterate function value gap bach_non-asymptotic_2011 ; taylor_stochastic_2019 ; liu_revisiting_2023 . At the time of writing this document, it was unclear whether SGD without stronger variance assumptions could enjoy last-iterate rates, hence justifying our choice of performance metric. During the review phase of the manuscript, two independent works have answered this question positively attia_fast_2025 ; garrigos_lastiterate_2025 , opening the door for the possibility to extend our results to the setting of last-iterate rates.
Paper Organization. This paper is organized as follows: We present the formal problem and the assumptions in Section 2. In Section 3, we outline the proof format of our main results. Sections 4 and 5 are devoted to our results (new bounds, most of which being bias-optimal) in the convex and strongly convex settings, respectively. In Section 6 we present the application of our results to the Stochastic Proximal algorithm. Section 7 discusses our PEP framework and the numerical optimality of our variance terms. The technical proofs are gathered in the appendices.
2 Preliminaries
We first formalize the problem statement presented in Problem ($\mathbb{E}$-minimization).
Problem 1(Main problem)
Let $\mathcal{H}$ be a real Hilbert space with associated inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$. Let $\{f_{i}\}_{i\in\mathcal{I}}$ be a family of real-valued functions $f_{i}\colon\mathcal{H}\to\mathbb{R}$, where $\mathcal{I}$ is a (possibly infinite) set of indices. We consider the problem of minimizing $f\coloneqq\mathbb{E}\mathopen{}\left[{f_{i}}\right]\mathclose{}$, where the expectation is taken over the indices $i\in\mathcal{I}$, with respect to some probability distribution $\mathcal{D}$ over $\mathcal{I}$. We assume that
@@ 2 Preliminaries
Within the context of Problem 1 the (SGD) algorithm is well defined. To analyze its performance, we will further assume the problem to be convex and smooth. We define a setting combining both convexity and strong convexity, which allows us to unify parts of the analyses.
{assumption}
[Smoothness and convexity] Considering Problem 1, we assume that
  • $\bullet$
@@
    there exists $L\in(0,+\infty)$ such that each $f_{i}$ is $L$-smooth, i.e. $\nabla f_{i}\colon\mathcal{H}\to\mathcal{H}$ is $L$-Lipschitz continuous;
  • $\bullet$
@@
    there exists $\mu\in[0,+\infty)$ such that each function $f_{i}$ is $\mu$-strongly convex. When $\mu=0$, this means that $f_{i}$ is convex.
@@
Smoothness and strong convexity can be characterized by the means of a variational inequality, which will be central to the derivation of our results.
Lemma 1(Expected cocoercivity)
Let Assumption 1 hold. Then, for every $(x,y)\in\mathcal{H}^{2}$,
@@
$\frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(y)-\nabla f_{i}(x)\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|y-x\|^{2}\leq(L-\mu)\left(f(y)-f(x)\right)+\langle\mu\nabla f(y)-L\nabla f(x),y-x\rangle.$ (EC)
Proof
@@
Because each $f_{i}$ is $L$-smooth and $\mu$-strongly convex, we can use (taylor_smooth_2017, , Theorem 4) and take an expectation over $i$.
We now turn to the constant $\sigma_{*}^{2}$ which controls the complexity of the SGD algorithm.
{assumption}
@@
[Finite solution variance] Considering Problem 1, we assume that the variance at the solution exists, namely that $\sigma^{2}_{*}\coloneqq\inf\limits_{x_{*}\in{\rm{argmin}}~f}\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]<+\infty$.
@@
We note that, under Assumption 1, the variance term $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$ does not depend on the choice of $x_{*}\in\operatorname{argmin}f$, as was already observed in (garrigos_handbook_2024, , Lemma 4.17).
Lemma 2
@@
Let Assumption 1 hold. Then $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$ is constant over ${\rm{argmin}}~f$.
Proof
@@
Let $x_{*},x_{*}^{\prime}$ be two minimizers of $f$. From Lemma 1 we know that (EC) holds true for the couple $(x_{*},x_{*}^{\prime})$, which means that
@@
$\frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\leq(L-\mu)\left(f(x_{*}^{\prime})-f(x_{*})\right)+\langle\mu\nabla f(x_{*}^{\prime})-L\nabla f(x_{*}^{\prime}),y-x\rangle.$
@@
Then, almost surely with respect to the distribution $\mathcal{D}$ over the indices $i\in\mathcal{I}$, we have $\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}=0$. This means that $\nabla f_{i}(x_{*}^{\prime})=\nabla f_{i}(x_{*})$ almost surely, from which we conclude that $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})\|^{2}}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$.
Assuming that the variance at the solution $\sigma_{*}^{2}$ exists is trivially satisfied in the case of the minimization of a finite sum.
Lemma 3
If Problem 1 treats a finite sum of functions, i.e. $\mathcal{I}$ is finite, then Assumption 2 holds true.
Even for true expectation-minimization problems, where $\mathcal{I}$ is infinite, Assumption 2 is very mild. In particular, this assumption is automatically verified in problems where the loss functions $f_{i}$ are nonnegative, which is standard for problems arising in inverse problems and machine learning.
Lemma 4
Consider Problem 1 and assume that the functions $f_{i}$ are $L$-smooth and are uniformly bounded from below functions: there exists $r\in\mathbb{R}$ such that $f_{i}(x)\geq r$ for every $i\in\mathcal{I}$ and $x\in\mathcal{H}$. Then Assumption 2 is true, with moreover the information that
@@
Proof
Since each $f_{i}$ is $L$-smooth, we can write (garrigos_handbook_2024, , Lemma 2.28), for all $x\in\mathcal{H}$,
@@
$\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\leq 2L(f(x_{*})-r)=2L(\min f-r)<+\infty,$
which concludes since $\sigma_{*}^{2}=\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$.
We now highlight the connection between the variance at the solution $\sigma^{2}_{*}$ with the notion of interpolation. The following facts can be found in garrigos_handbook_2024 , which we have extended here from a finite-sum problem to our expectation-minimization Problem 1.
Definition 1(Interpolation)
Consider Problem 1. We say that interpolation holds for the family of functions $\{f_{i}\}_{i\in\mathcal{I}}$ if there exists $x_{*}$ which belongs to ${\rm{argmin}}~f_{i}$ for $\mathcal{D}$-almost every $i$.
Lemma 5
In the context of Problem 1, if interpolation holds, then $\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}=\inf f$ and $\cap_{i\in\mathcal{I}}\ {\rm{argmin}}~f_{i}={\rm{argmin}}~f$, where the intersection is to be understood for $\mathcal{D}$-almost every $i\in\mathcal{I}$.
Proof
@@
Since $f_{i}(x)-\inf f_{i}\geq 0$ for all $i$, it must hold that $f_{i}(x)-\inf f_{i}=0$ for almost every $i\in\mathcal{I}$, which proves the claim.
Lemma 6
In the context of Problem 1, if interpolation holds, then $\sigma_{*}^{2}=0$. This becomes an equivalence if all the functions $f_{i}$ are convex.
Proof